Information: In September of 2016 it was reported that AAA rated corporate bonds with 20 years to maturity were yielding 3.18%, while AAA rated municipal bonds of the same maturity yielded 2.57%


Question: Based on this information what was the implied marginal tax bracket (break-even tax bracket) at the time for bonds of this maturity and risk level? If a County Water Reclamation District revenue bond yields 2.7%, what is its tax-equivalent yield for investors in the 28% tax bracket?
Answer #1

First question,

Corporate Bonds maturity yielded = 3.18%

Municipal bonds maturity yielded = 2.57%

Putting the given value in the expression we get,

(1 - t)0.0318 = 0.0257

t = 1 - (0.0257/0.0318)

t = 19.18%

Therefore, implied marginal tax rate = 19.18%

Second question,

Given,

bond yields = 2.7%

tax bracket = 28%

Therefore, putting value in the expression we get,

Tax Equivalent Yield = 0.027/(1 - 0.28) = 3.75%

Please don't forget to rate the answer if its helpful, thank you.